The perimeter of a rectangle is 62m. If the width was doubled and the length was increased by 21m, the perimeter would be 120m.
What is the length and width of the rectangle?

Answer by uday1435(57)

Let the length and width be l and w respectively
Perimeter = 2( l + w ) = 62 (1)
Also given 2(l + 21 + 2w ) = 120 (2)
Ie 2(l + w) + 2(21 + w) = 120
62 + 42 + 2w = 120
104 + 2w = 120
2w = 16, so w = 8
Plugging in the value of w in eqn(1)
2(l + 8) = 62
2l + 16 = 62
2l = 62 – 16 = 46 so l = 46/2 = 23
Ans: Length = 23 meters and Width = 8 meters

If you still require further clarifications please write to
udayakumar.t.r@gmail.com